OpenStudy (followthefez):

Can someone please explain how to do unit conversions? I understand how to convert things when nothing is squared. But I get confused when something is squared.

OpenStudy (followthefez):

I understand how to convert metres to feet, and meters per second to kilometers per minute etc etc. What I get confused with is when a question asks to convert something from \[dollars / m^{2}\] to\[dollars/yard^{2}\] or from something like mL to L^2 etc. I am never too sure what I should be doing with the squared sign! Someone explain please?

OpenStudy (unklerhaukus):

for example ; convert \(1[\text{cm}]^3\) to \([\text{m}]^3\)

OpenStudy (unklerhaukus):

\[1[\text{cm}]^3=\big(10^{-2}[\text{m}\big])^3=10^{-6}[\text{m}]^3\]

OpenStudy (unklerhaukus):

\[6\,[$/\text{yd}^2]=6\,[$/\text{yd}^2]\times\big(1\big)^2\] \[=6\,[$/\text{yd}^2]\times\big(1.09[\text{yd}/\text m]\big)^2\] \[=6\,[$/\text{yd}^2]\times1.09^2[\text{yd}^2/\text m^2]\] \[=6\times1.09^2\,[$/\text m^2]\]